OpenStudy (anonymous):

How do I find the domain and range for f(x) = (1/2)e^x , when f(x)=0?

OpenStudy (anonymous):

you can take \(e\) to any power you like, so the domain would be all real numbers

OpenStudy (anonymous):

So is the range also all real numbers because of that?

OpenStudy (anonymous):

Wait i forgot to mention the question is "find the domain and range of the following, when f(x)=0"

OpenStudy (anonymous):

that question makes no sense

OpenStudy (anonymous):

the range is not all real numbers although \(x\) can be any number, \(e^x>0\) no matter what so the range is \((0,\infty)\)

OpenStudy (anonymous):

Okay to be clear "find the domain and range for the following. Find when f(x)=0" then it gives me f(x)=(1/2)e^x

OpenStudy (anonymous):

ok so that is two different questions well actually 3 1) domain 2) range 3) solve \(\frac{1}{2}e^x=0\)

OpenStudy (anonymous):

1) domain is all real numbers, because you can raise \(e\) to any power you like 2) range is \((0,\infty)\) since \(e^x>0\) for all \(x\) 3) no solution for the reason given in answer 2

OpenStudy (anonymous):

okay thanks for the help